# Complexity of Weighted First-Order Model Counting in the Two-Variable Fragment with Counting Quantifiers: A Bound to Beat

### v1

@@ 4.1 Canonical Models
Now, let us transform $\omega$ into $\omega_{t}$ which will be another model of $\Gamma^{*}$.
First, we separate all atoms in $\omega$ (atoms true in $\omega$) without the predicates $f_{i}/2$ and $A_{i}/1$ into the set $\mathcal{R}_{0}$, and atoms from $\omega\setminus\mathcal{R}_{0}$ not containing the constant $t$ into $\mathcal{R}_{1}$. Second, we define an auxiliary injective function $g:\mathcal{A}^{t}\mapsto\left[k\right]$ mapping elements of $\mathcal{A}^{t}$ to the first $|\mathcal{A}^{t}|$ positive integers, i.e.,
$g(j)=|\{j^{\prime}\in\mathcal{A}^{t}\;|\;j^{\prime}\leq j\}|.$
Third, we define a set of atoms $\mathcal{A}^{new}$ such that
$\mathcal{A}^{new}=\{A_{{g(j)}}(t)\;|\;\omega\models A_{j}(t)\},$
i.e., we accumulate the Skolemization atoms with the constant $t$ that are satisfied in $\omega$ and we change their indices to the first $|\mathcal{A}^{t}|$ positive integers. Next, we do a similar thing for atoms with $f_{i}$’s. Note, that we use the same function $g$ as for $\mathcal{A}^{t}$. Hence, we construct a set
$\mathcal{F}^{new}=\{f_{{g(j)}}(t,t^{\prime})\;|\;\omega\models f_{j}(t,t^{\prime}),t^{\prime}\in\Delta\}.$
@@
$\omega_{t}=\mathcal{R}_{0}\cup\mathcal{R}_{1}\cup\mathcal{A}^{new}\cup\mathcal{F}^{new}.$
@@
To prove the claim, it is sufficient to note that we can permute the indices $i$ and the sentence $\Gamma^{*}$ will remain the same. The transformation simply permutes the indices so that the satisfied atoms with predicates $A_{i}$ and the constant $t$ have the lowest possible indices. ∎
Now, suppose that we take some model of $\Gamma^{*}$ and we repeatedly perform the transformation described above for all domain elements, i.e., for some ordering of the domain such as $\Delta=\{t_{1},t_{2},\ldots,t_{n}\}$, we construct $\omega_{t_{1}}$ from $\omega$, then we construct $\omega_{t_{2}}$ from $\omega_{t_{1}}$ etc., until we obtain $\omega_{t_{n}}=\omega^{*}$. Note that there can be several models $\omega$ that lead to the same $\omega^{*}$, thus $\omega^{*}$ effectively induces an equivalence class.
@@
It follows from the proof of Lemma 7 that the atoms of $A_{i}$’s that are true in $\omega^{*}$ have the lowest possible indices. Hence, for a fixed $t\in\Delta$, the number of models that lead to $\omega^{*}$ depends only on the number of ways that we can split the indices between the satisfied and unsatisfied atoms. There are $k$ indices to choose from and, for a fixed assignment, the set $\mathcal{A}^{t}$ holds the indices of satisfied $A_{j}(t)$’s, which leads directly to Equation 26.
@@ 4.2 The New Encoding
One more thing to consider are weights for the predicates $C_{i}$. A particular atom $C_{j}(t)$ is satisfied in a model $\omega$, if the model is a canonical model corresponding to $C_{j}(t)$ according to the sentence $\Gamma_{C}$. Following Lemma 8, such model represents an entire set of A-equivalent models, each with the same weight. Hence, if $C_{j}(t)$ is satisfied, we should count the model weight of $\omega$ as many times as how many A-equivalent models to $\omega$ there are. If $C_{j}(t)$ is, on the other hand, unsatisfied, we do not want to change the weight in any way. Therefore, we set $w(C_{i})=\binom{k}{i}$ and $\overline{w}(C_{i})=1$ for all $i\in\left[k\right]$.
@@
For any sentence $\Gamma\in\textbf{C}^{2}{}$ such as the one in Equation 25 and $\Gamma^{new}\in\textbf{FO}^{2}+\textbf{CC}{}$ obtained from $\Gamma$ as in Equation 27, for any $n\in\mathds{N}$ and any weights $(w,\overline{w})$ extended for predicates $C_{i}$ as above, it holds that
@@
$\Gamma_{C}$ encodes each canonical model of $\Gamma^{*}$ using a single $C_{j}$ predicate indicating that exactly $A_{1}(t),A_{2}(t),\ldots,A_{j}(t)$ are satisfied for any $t\in\Delta$. For all $i\in\{0,1,\ldots,k\}$, $\Gamma_{C}$ allows at most one $C_{i}(t)$ to be true for any $t$. The disjuction in $\Gamma^{new}$ requires at least one $C_{i}(t)$ to be true. Combined, each model of $\Gamma^{new}$ is one of canonical models of $\Gamma^{*}$. Hence, we are only counting canonical models of $\Gamma^{*}$.
It follows from Lemma 8 that for a canonical model $\omega$ represented by $C_{i}$, there are $\binom{k}{i}$ models of $\Gamma^{*}$ that are A-equivalent to $\omega$. Hence, when we have one canonical model, it represents $\binom{k}{i}$ models that must all be considered to obtain the correct weighted model count leading to the weight $w(C_{i})=\binom{k}{i}$ for each $C_{i}$.
So far, we have shown that we are counting models of $\Gamma^{*}$. However, we must still deal with the over counting introduced by using Lemma 4 to turn $\Gamma$ into $\Gamma^{*}$. Therefore, we use the same factor as in Lemma 4 to only count models of $\Gamma$. ∎
@@ 4.3 The Improved Upper Bound
Consider an arbitrary $\textbf{C}^{2}{}$ sentence rewritten as
where $\Gamma\in\textbf{FO}^{2}+\textbf{CC}$. For any $n\in\mathds{N}$ and any fixed weights $(w,\overline{w})$, $\textup{WFOMC}(\varphi,n,w,\overline{w})$ can be computed in time $\mathcal{O}(n^{\alpha^{\prime}}\cdot n^{1+\prod_{i=1}^{m}\beta\cdot\gamma^{\prime}(k_{i})\cdot p})$, where $p$ is the number of valid cells of $\Gamma$ and $\gamma^{\prime}(k)=\mathcal{O}(k^{2}+2k+1)$.
@@
Most of the proof is identical to the proof of Theorem 2. What remains is to derive $\gamma^{\prime}$.
@@
As we already know, sentence $\Gamma$ causes the largest increase in the number of valid cells. With the old encoding, most of truth values of the Skolemization atoms with predicates $A_{i}$ were unconstrained leading to an exponential blowup. Due to Equation 27, that is no longer the case. The sentence $\Gamma_{C}$ forces $C_{j}(t)$ to be true if and only if $A_{1}(t),A_{2}(t),\ldots,A_{j}(t)$ are true and all other $A_{i}(t)$ with $j<i\leq k$ are false. Hence, if $C_{j^{\prime}}(t)$ is true, then all other $C_{j}(t)$ with $j^{\prime}\neq j$ are false. Moreover, at least one $C_{i}(t)$ must be true due to the final disjunction in Equation 27.
Therefore, we only have $(k+1)$ possibilities, as how to assign truth values to atoms with $C_{i}$ predicates. The truth values of atoms with $A_{i}$’s are then directly determined without any more degrees of freedom. Hence, for both cases considered in the proof of Lemma 6, we receive a factor $(k+1)$ instead of the exponential. Thus, we obtain $\gamma^{\prime}=\mathcal{O}((k+1)p+k(k+1)p)\in\mathcal{O}((k^{2}+2k+1)p)$. ∎
@@ 5 Experiments
In this section, we support our theoretical findings by providing time measurements for various WFOMC computations using both the old and the new $\textbf{C}^{2}$ encoding. We also provide tables comparing the number of valid cells $p$, which determines the polynomial degree, giving a more concrete idea on the speed-up provided by the new encoding.
For all of our experiments, we used FastWFOMC.jl, [^5] https://github.com/jan-toth/FastWFOMC.jl an open source Julia implementation of the FastWFOMC algorithm (van Bremen and Kuželka, 2021a), which is arguably the state-of-the-art, reported by its authors to outperform the first approach to computing WFOMC in a lifted manner, i.e, ForcLIFT, [^6] https://dtaid.cs.kuleuven.be/wfomc which is based on knowledge compilation (Van den Broeck et al., 2011).
Apart from time measurements for $\textbf{C}^{2}$ sentences, we also inspect sentences from one of the domain-liftable $\textbf{C}^{2}$ extensions, namely $\textbf{C}^{2}$ with the linear order axiom (Tóth and Kuželka, 2023). Apart from time measurements on sentences from that fragment, we also use it to encode a Markov Logic Network and perform exact lifted inference therein.
Most of our experiments were performed in a single thread on a computer with an AMD Ryzen 5 7500F CPU running at speeds 3.4GHz and 32 GB RAM. Problems containing the linear order axiom, which have considerably higher memory requirements, were solved using a machine with AMD EPYC 7742 CPU running at speeds 2.25GHz and 512 GB of RAM.
5.1 Performace Measurements
5.1.1 Sentences with Counting Quantifiers
@@
$\Gamma_{1}$ $=\left(\forall x\;\neg E(x,x)\right)\wedge\left(\forall x\forall y\;E(x,y)\Rightarrow E(y,x)\right)\wedge\left(\forall x\exists^{=k}y\;E(x,y)\right).$
In our experiments, we simply count the number of such graphs on $n$ vertices, i.e., we set $w(E)=\overline{w}(E)=1$ and we compute $\textup{WFOMC}(\Gamma_{1},n,w,\overline{w})$ for gradually increasing domain sizes $n$. Since our new encoding is aimed at improving the runtime with respect to the counting parameter $k$, we do the computation for several different parameters $k$ as well.
Figure 1 shows the measured runtimes for $k\in\{3,4,5\}$. As one can observe, the new encoding surpasses the old in each case. The difference may not seem as distinct for $k=3$ compared to $k=4$ or $k=5$ but note that it is still substantial. For one, runtime for $n=51$ already exceeded runtime of $1000$ seconds in case of the old encoding, whereas the new encoding did not reach that value even for $n=70$. Additionally to the figure, Table 1 shows the number of valid cells $p_{old}$ produced by the old encoding of $\Gamma_{1}$ into $\textbf{FO}^{2}+\textbf{CC}$ and $p_{new}$ produced by the new one. As one can observe, e.g., for $k=5$, the new encoding reduces the runtime from $\mathcal{O}(n^{33})$ to $\mathcal{O}(n^{7})$.
@@
Next, we can extend the $k$-regular graph definition to also include a coloring by $l$ colors, effectively obtaining a logical description of $k$-regular $l$-colored graphs. [^7] Note that $l$-colored graphs are graphs along with a coloring using at most $l$ colors. That is different from $l$-colorable graphs which is a set of graphs that can be colored using $l$ colors. Sentence $\Gamma_{2}$ encodes such graphs and we again count the number of such graphs on larger and larger domains for various values $k$ and $l$ using WFOMC.
$\Gamma_{2}$ $=\forall x\;\neg E(x,x)\wedge\forall x\forall y\;E(x,y)\Rightarrow E(y,x)$
$\wedge\bigwedge_{1\leq i<j\leq l}\left(\forall x\;C_{i}(x)\Rightarrow\neg C_{j}(x)\right)$
@@
$\wedge\forall x\exists^{=k}y\;E(x,y)\wedge\forall x\;\bigvee_{i=1}^{l}C_{i}(x)$
The runtime measurements are available in Figure 2. Similarly to the previous case, Table 2 shows number of valid cells for various scenarios, even for some that are not included in Figure 2 due to their computational demands.
@@
In the examples so far, our $\textbf{C}^{2}$ sentences contained only one $\forall\exists^{=k}$ sentence. Directed $k$-(bi)regular graphs without loops, requiring both the number of outgoing and incoming edges to be exactly $k$ for each vertex, contain two such counting sentences, putting larger emphasis on our upper bounds. We encode such graphs in $\Gamma_{3}$:
$\Gamma_{3}$ $=\left(\forall x\;\neg E(x,x)\right)\wedge\left(\forall x\exists^{=k}y\;E(x,y)\right)\wedge\left(\forall x\exists^{=k}y\;E(y,x)\right)$
Runtimes for counting over such digraphs may be inspected in Figure 3. Table 3 depicts valid cells counts, which turn out to be quite high. Due to those high values, which end up being polynomial degrees in our upper bounds, the problem is considerably more difficult. Even for $k=3$, the old encoding was extremely inefficient, hence, we only provide runtimes for counting with the new encoding in that case. [^8] Note, that for the case of $k=3$, the corresponding OEIS sequence at https://oeis.org/A007105 only contains elements up to $n=14$, which is the same as we managed to compute within a time limit of $10^{5}$ seconds.

### v2

@@ 4.1 Canonical Models
Now, let us transform $\omega$ into $\omega_{t}$, which will be another model of $\Gamma^{*}$.
First, we separate all atoms in $\omega$ (atoms true in $\omega$) without the predicates $f_{i}/2$ and $A_{i}/1$ into the set $\mathcal{R}_{0}$, atoms on $A_{i}/1$ not containing the constant $t$ into $\mathcal{R}_{t}^{A}$ and atoms on $f_{i}/2$ not containing the constant $t$ on the first position into $\mathcal{R}_{t}^{f}$.
Second, we define an auxiliary injective function $g_{t}:\mathcal{A}^{t}\mapsto\left[k\right]$ mapping elements of $\mathcal{A}^{t}$ to the first $|\mathcal{A}^{t}|$ positive integers, i.e.,
$g_{t}(j)=|\{j^{\prime}\in\mathcal{A}^{t}\;|\;j^{\prime}\leq j\}|$ (26)
Third, we define a set of atoms $\mathcal{A}^{new}_{t}$ such that
$\mathcal{A}^{new}_{t}=\{A_{{g_{t}(j)}}(t)\;|\;\omega\models A_{j}(t)\},$
i.e., we accumulate the Skolemization atoms with the constant $t$ that are satisfied in $\omega$ and we change their indices to the first $|\mathcal{A}^{t}|$ positive integers. Next, we do a similar thing for atoms with $f_{i}$’s and the constant $t$ at the first position. Note that we use the same function $g_{t}$ that was defined (with respect to $\mathcal{A}^{t}$) in Equation 26. Hence, we construct a set
$\mathcal{F}^{new}_{t}=\{f_{{g_{t}(j)}}(t,t^{\prime})\;|\;\omega\models f_{j}(t,t^{\prime}),t^{\prime}\in\Delta\}.$
@@
$\omega_{t}=\mathcal{R}_{0}\cup\mathcal{A}^{new}_{t}\cup\mathcal{F}^{new}_{t}\cup\bigcup_{t^{\prime}\in\Delta\setminus\{t\}}\left(\mathcal{R}_{t^{\prime}}^{A}\cup\mathcal{R}_{t^{\prime}}^{f}\right).$
@@
To prove the claim, it is sufficient to note that we can permute the indices $i$, and the sentence $\Gamma^{*}$ will remain the same. The transformation permutes the indices so that the atoms $A_{i}(t)$ that are satisfied have the lowest possible indices. ∎
Now, suppose that we take some model of $\Gamma^{*}$ and we repeatedly perform the transformation described above for all domain elements, i.e., for some ordering of the domain such as $\Delta=\{t_{1},t_{2},\ldots,t_{n}\}$, we construct $\omega_{t_{1}}$ from $\omega$, then we construct $\omega_{t_{2}}$ from $\omega_{t_{1}}$ and so on, until we obtain $\omega_{t_{n}}=\omega^{*}$. Note that several models $\omega$ can lead to the same $\omega^{*}$. Thus, $\omega^{*}$ effectively induces an equivalence class.
@@
It follows from the proof of Lemma 7 that the atoms of $A_{i}$’s that are true in $\omega^{*}$ have the lowest possible indices. Hence, for a fixed $t\in\Delta$, the number of models that lead to $\omega^{*}$ depends only on the number of ways that we can split the indices between the satisfied and the unsatisfied atoms. There are $k$ indices to choose from, and for a fixed assignment, the set $\mathcal{A}^{t}$ holds the indices of satisfied $A_{j}(t)$’s, which leads directly to Equation 27.
@@ 4.2 The New Encoding
One more thing to consider is the weights for the predicates $C_{i}$. A particular atom $C_{j}(t)$ is satisfied in a model $\omega$ if the model is a canonical model corresponding to $C_{j}(t)$ according to the sentence $\Gamma_{C}$. Following Lemma 8, such a model represents an entire set of A-equivalent models, each with the same weight. Hence, if $C_{j}(t)$ is satisfied, we should count the model weight of $\omega$ as many times as how many A-equivalent models to $\omega$ there are. If $C_{j}(t)$ is, on the other hand, unsatisfied, we want to keep the weight the same. Therefore, we set $w(C_{i})=\binom{k}{i}$ and $\overline{w}(C_{i})=1$ for all $i\in\left[k\right]$.
@@
For any sentence $\Gamma\in\textbf{C}^{2}{}$ such as the one in Equation 25 and $\Gamma^{new}\in\textbf{UFO}^{2}+\textbf{CC}{}$ obtained from $\Gamma$ as in Equation 28, for any $n\in\mathds{N}$ and any weights $(w,\overline{w})$ extended for predicates $C_{i}$ as above, it holds that
@@
$\Gamma_{C}$ encodes each canonical model of $\Gamma^{*}$ using a single $C_{j}$ predicate indicating that for any $t\in\Delta$, $A_{1}(t),A_{2}(t),\ldots,A_{j}(t)$ are satisfied. For all $i\in\{0,1,\ldots,k\}$, $\Gamma_{C}$ allows at most one $C_{i}(t)$ to be true for any $t$. The disjunction in $\Gamma^{new}$ requires at least one $C_{i}(t)$ to be true. Combined, each model of $\Gamma^{new}$ is one of the canonical models of $\Gamma^{*}$. Hence, we are only counting canonical models of $\Gamma^{*}$.
It follows from Lemma 8 that for a canonical model $\omega$ represented by $C_{i}$, there are $\binom{k}{i}$ models of $\Gamma^{*}$ that are A-equivalent to $\omega$. Hence, when we have one canonical model, it represents $\binom{k}{i}$ models that must all be considered to obtain the correct weighted model count, leading to the weight $w(C_{i})=\binom{k}{i}$ for each $C_{i}$.
So far, we have shown that we are counting models of $\Gamma^{*}$. However, we must still deal with the overcounting introduced by applying Lemma 4 to turn $\Gamma$ into $\Gamma^{*}$. Therefore, we use the same factor as in Lemma 4 to only count models of $\Gamma$. ∎
@@ 4.3 The Improved Upper Bound
Consider an arbitrary $\textbf{C}^{2}{}$ sentence rewritten as $\varphi=\Gamma\wedge\Phi$, where $\Gamma\in\textbf{FO}^{2}+\textbf{CC}$ and
$\Phi=\bigwedge_{i=1}^{m}\left(\forall x\;P_{i}(x)\Leftrightarrow\left(\exists^{=k_{i}}y\;R_{i}(x,y)\right)\right).$
For any $n\in\mathds{N}$ and any fixed weights $(w,\overline{w})$, $\textup{WFOMC}(\varphi,n,w,\overline{w})$ can be computed in time
$\mathcal{O}(n^{\alpha^{\prime}}\cdot n^{1+p\cdot\prod_{i=1}^{m}\beta\cdot\gamma^{\prime}(k_{i})}),$
where $p$ is the number of valid cells of $\Gamma$ and $\gamma^{\prime}(k)=\mathcal{O}(k^{2}+2k+1)$.
@@
Let us derive $\gamma^{\prime}$. All other values can be derived identically as in the proof of Theorem 2.
@@
As we already know, sentence $\Gamma$ causes the largest increase in the number of valid cells.
With the old encoding, most of the truth values of the Skolemization atoms with predicates $A_{i}$ were unconstrained, leading to an exponential blowup. Due to Equation 28, that is no longer the case. The sentence $\Gamma_{C}$ forces $C_{j}(t)$ to be true if and only if $A_{1}(t),A_{2}(t),\ldots,A_{j}(t)$ are true and all other $A_{i}(t)$ with $j<i\leq k$ are false. Hence, if $C_{j^{\prime}}(t)$ is true, then all other $C_{j}(t)$ with $j^{\prime}\neq j$ are false. Moreover, at least one $C_{i}(t)$ must be true due to the final disjunction in Equation 28.
Therefore, we only have $(k+1)$ possibilities for assigning truth values to atoms with $C_{i}$ predicates. The truth values of atoms with $A_{i}$’s are then directly determined without any more degrees of freedom. Hence, for both cases considered in the proof of Lemma 6, we receive a factor $(k+1)$ instead of the exponential. Thus, we obtain
$\gamma^{\prime}=\mathcal{O}(p\cdot(k+1)+p\cdot k\cdot(k+1))\in\mathcal{O}(p\cdot(k^{2}+2k+1)).$
@@ 5 Experiments
In this section, we support our theoretical findings by providing time measurements for various WFOMC computations using both the old and the new $\textbf{C}^{2}$ encoding. We also provide tables comparing the number of valid cells $p$, which determines the polynomial degree, giving a more concrete idea of the speedup provided by the new encoding. Last, but not least, we provide an experiment performing lifted inference over a Markov Logic Network (Richardson and Domingos, 2006) defined using the language of $\textbf{C}^{2}$ extended with the linear order axiom.
For all of our experiments, we used FastWFOMC.jl, [^6] https://github.com/jan-toth/FastWFOMC.jl an open source Julia implementation of the FastWFOMC algorithm (van Bremen and Kuželka, 2021a), which is arguably the state-of-the-art, reported by its authors to outperform the first approach to computing WFOMC in a lifted manner, i.e, ForcLIFT, [^7] https://dtaid.cs.kuleuven.be/wfomc which is based on knowledge compilation (Van den Broeck et al., 2011). Apart from time measurements for $\textbf{C}^{2}$ sentences, we also inspect sentences from one of the domain-liftable $\textbf{C}^{2}$ extensions, namely $\textbf{C}^{2}$ with the linear order axiom (Tóth and Kuželka, 2023).
Most of our experiments were performed in a single thread on a computer with an AMD Ryzen 5 7500F CPU running at 3.4GHz and having 32 GB RAM. Problems containing the linear order axiom, which have considerably higher memory requirements, were solved using a machine with AMD EPYC 7742 CPU running at 2.25GHz and having 512 GB of RAM.
5.1 Performance Measurements
In this section, we present observed running times for several problems specified using the language of $\textbf{C}^{2}$ or its extension by the linear order axiom. Most of our experiments show simply counting the number of graphs with some specified properties. The advantage of using counting problems is that we can easily check the obtained results using the On-Line Encyclopedia of Integer Sequences (OEIS Foundation Inc., 2023). To perform counting as opposed to weighted counting, we simply set the weights (both positive and negative) of all predicates in the input formula to one.
5.1.1 Counting $k$-regular Graphs
@@
$\Gamma_{1}$ $=\left(\forall x\;\neg E(x,x)\right)\wedge\left(\forall x\forall y\;E(x,y)\Rightarrow E(y,x)\right)$
$\wedge\left(\forall x\exists^{=k}y\;E(x,y)\right).$
Figure 1 shows the measured running times of $\textup{WFOMC}(\Gamma_{1},n,1,1)$ for $k\in\{3,4,5\}$. As one can observe, the new encoding surpasses the old in each case. The difference may not seem as distinct for $k=3$ compared to $k=4$ or $k=5$, but it is still substantial. For one, runtime for $n=51$ already exceeded runtime of $1000$ seconds in the case of the old encoding, whereas the new encoding did not reach that value even for $n=70$. Additionally to the figure, Table 1 shows the number of valid cells $p_{old}$ produced by the old encoding of $\Gamma_{1}$ into $\textbf{UFO}^{2}+\textbf{CC}$ and $p_{new}$ produced by the new one. As one can observe, e.g., for $k=5$, the new encoding reduces the runtime from $\mathcal{O}(n^{33})$ to $\mathcal{O}(n^{7})$.
@@
5.1.2 Counting $k$-regular $l$-colored Graphs
Next, we can extend the $k$-regular graph definition to also include a coloring by $l$ colors, effectively obtaining a logical description of $k$-regular $l$-colored graphs. [^8] Note that $l$-colored graphs are graphs along with a coloring using at most $l$ colors. That is different from $l$-colorable graphs which is a set of graphs that can be colored using $l$ colors. Sentence $\Gamma_{2}$ encodes such graphs.
$\Gamma_{2}$ $=\left(\forall x\;\neg E(x,x)\right)\wedge\left(\forall x\forall y\;E(x,y)\Rightarrow E(y,x)\right)$
$\wedge\left(\forall x\exists^{=k}y\;E(x,y)\right)$
$\wedge\left(\forall x\;\left(\bigvee_{i=1}^{l}C_{i}(x)\right)\right)\wedge\bigwedge_{1\leq i<j\leq l}\left(\forall x\;C_{i}(x)\Rightarrow\neg C_{j}(x)\right)$
@@ 5.1.2 Counting $k$-regular $l$-colored Graphs
The runtime measurements of $\textup{WFOMC}(\Gamma_{2},n,1,1)$ are available in Figure 2. Similarly to the previous case, Table 2 shows the number of valid cells for various scenarios, even for some that are not included in Figure 2 due to their computational demands.
@@
5.1.3 Counting $k$-regular Directed Graphs
In the examples so far, our $\textbf{C}^{2}$ sentences contained only one $\forall\exists^{=k}$-quantified sentence. Let us now see a case with more than one counting sentence, which should put an even larger emphasis on our upper bounds. Directed $k$-(bi)regular graphs without loops require both the number of outgoing and incoming edges to be exactly $k$ for each vertex. We can encode them using $\Gamma_{3}$, which uses two counting sentences:
$\Gamma_{3}$ $=\left(\forall x\;\neg E(x,x)\right)$
$\wedge\left(\forall x\exists^{=k}y\;E(x,y)\right)$
$\wedge\left(\forall x\exists^{=k}y\;E(y,x)\right)$
Runtimes for counting over such digraphs may be inspected in Figure 3. Table 3 depicts valid cell counts, which turn out to be quite high. Due to those high values, which end up being polynomial degrees in our upper bounds, the problem is considerably more difficult. Even for $k=3$, the old encoding was extremely inefficient. Hence, we only provide runtimes for counting with the new encoding in that case. [^9] Note that for the case of $k=3$, the corresponding OEIS sequence at https://oeis.org/A007105 also contains only elements up to $n=14$, which is the same as we managed to compute within a time limit of $10^{5}$ seconds.